When a heating coil is immersed in melting ice, the ice melts at a rate of 10grams per minute. If the current in the coil is doubled while its resistance remains constant, the mass of ice that will melt per minute is?



1
Expert's answer
2021-11-05T18:19:33-0400

"Q_1=\lambda m,"

"Q_2=I^2Rt,"

"Q_1=Q_2,"

"\lambda m=I^2Rt,"

"m\sim I^2,"

"m_2=m_1\frac{(2I)^2}{I^2}=4m_1=40~g."